# Adversarial Reinforcement Learning Framework for ESP Cheater Simulation

### v1

@@ 1 Introduction
Cheating is a persistent issue in digital games that significantly violates the fairness and degrades the user experience. If cheating is not properly addressed, it may drive players away from the game, ultimately leading to reduced revenue for game developers. It has been reported to cause an estimated loss of \$29 billion and to drive away 78% of gamers for a year (Irdeto, 2022).
One of the most representative forms of cheating is Extra-Sensory Perception (ESP), which allows cheaters to access hidden game information. A common example is the use of wallhacks in first-person shooter (FPS) games: while normal players cannot see opponents behind walls, cheaters can, enabling them to avoid unexpected attacks and easily eliminate enemies. These unfair advantages severely disrupt the balance of the game, especially in settings where players must make decisions and formulate strategies based on incomplete information. ESP cheats are notoriously hard to prevent due to their passive nature: they do not modify game files or client-side data. In addition, their usage is often indistinguishable from normal gameplay when observed from a third-person perspective, making them especially challenging to detect through in-game monitoring or user reports (PUBG: BATTLEGROUNDS Anti-Cheat Team, 2024).
@@
Next, cheaters continuously evolve their strategies to evade detection (Tao et al., 2018; Jonnalagadda et al., 2021). When a cheat detector is developed and deployed, cheaters adapt by limiting their cheat usage just enough to avoid detection. For example, if a simple threshold-based detection algorithm is used, cheaters will intentionally operate just below the threshold to remain undetected. As a result, the statistical characteristics of past cheater data differ from those of recent data. Since only recent data is useful for developing effective detectors, it becomes difficult to collect a large amount of high-quality training and test data.
@@
To verify the effectiveness of the framework, we implement Gridworld and Blackjack environments that can simulate both fully observable and partially observable agents. We also use various designs of the detector based on the trajectory, the trajectory length, and the reward. With these diverse configurations, we conduct extensive experiments, demonstrating the effectiveness of our framework in modeling adaptive cheaters. Our code is available at https://github.com/yunik1004/esp-simulator.
@@ 2 Related Works
Several studies (Alayed et al., 2013; Tao et al., 2018; Jonnalagadda et al., 2021; Pinto et al., 2021; Kanervisto et al., 2022; Zhang et al., 2024) have explored cheat detection in video games, proposing a variety of approaches for identifying different types of malicious behaviors. Alayed et al. (2013) used a support vector machine (SVM) (Suthaharan, 2016) to detect aimbots based on server-side logs from online FPS games. Tao et al. (2018) proposed NGUARD, a bot detection framework for multiplayer online role-playing games (MMORPGs), which combines supervised and unsupervised learning to detect known bots and discover previously unseen ones. Jonnalagadda et al. (2021) introduced a vision-based deep neural network (DNN) detector to identify illicit on-screen overlays in Counter-Strike: Global Offensive (CS:GO) (Valve, 2012). In addition, they applied the interval bound propagation method (Gowal et al., 2018) to defend against adversarial attacks targeting the detection system. Pinto et al. (2021) formulated sequences of keyboard and mouse events as a multivariate time series, and employed a convolutional neural network (CNN) (O’shea & Nash, 2015) to detect triggerbots and aimbots in CS:GO. Kanervisto et al. (2022) developed an aimbot that mimics human behavior using a generative adversarial network (GAN) (Goodfellow et al., 2020), and evaluated the performance of multiple detection methods against it. Zhang et al. (2024) proposed HAWK, an anti-cheat framework for CS:GO, which leverages machine learning to mimic the decision process of human experts in detecting wallhacks and aimbots. Despite recent advances, most approaches assume that cheaters behave in fixed, non-adaptive ways. It overlooks the fact that real-world cheaters often adjust their strategies to avoid detection, highlighting the need for simulation frameworks that can model such adaptive behaviors.
@@
Adversarial learning traces its roots to the GAN (Goodfellow et al., 2020), where a generator and a discriminator engage in a minimax game to match data distributions. Ho & Ermon (2016) later carried this game-theoretic structure over to reinforcement learning with generative adversarial imitation learning (GAIL), which trains a policy to fool a discriminator that separates expert and learner trajectories, thereby reproducing the expert’s state–action occupancy without an explicit reward-recovery step. For robustness, Pinto et al. (2017) proposed robust adversarial reinforcement learning (RARL), where an adversary injects disturbance forces into the simulator so that the resulting policy remains stable under variations in friction, mass, and other model mismatches. Dennis et al. (2020) later introduced protagonist–antagonist induced regret environment design (PAIRED), training an environment designer and an antagonist demonstrator to build a curriculum of tasks that are solvable yet currently unsolved by the protagonist, thereby boosting zero-shot transfer. Although these lines of work pursue diverse goals in imitation (Ho & Ermon, 2016), robustness (Pinto et al., 2017), curriculum-driven generalization (Dennis et al., 2020), and domain expansion (Cheng et al., 2025), they all leverage adversarial interaction to enhance policy capability. In contrast, our framework jointly trains an ESP cheater and a trajectory-level detector in a partially observable setting, exploiting dynamic cheating behavior that balances reward optimization and detection evasion.
@@ 3 ESP Simulation Framework
We model the game environment as a partially observable Markov decision process (POMDP) (Åström, 1965; Kaelbling et al., 1998), defined by the tuple $\langle\mathcal{S},\mathcal{A},T,r,\Omega,O,\gamma\rangle$, where $\mathcal{S}$ is the set of environment states, $\mathcal{A}$ is the set of actions, $T$ denotes the state transition probability distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\Omega$ is the set of observations available to the non-cheater player, $O$ is the observation function, defining the conditional observation probabilities, and $\gamma\in[0,1]$ is a discount factor. Consistent with previous works related to the trajectory feedback (Liu et al., 2019; Efroni et al., 2021), we assume $\gamma=1$ throughout this paper.
@@
    Given a trajectory, a sequence of states and actions $\tau=(s_{1},a_{1},s_{2},a_{2},\cdots,s_{t},a_{t})$, the detector estimates the probability that it was generated by the cheater. During the inference, the detector receives no information about the ground-truth label of the trajectory. On the other hand, it is allowed to access the ground-truth label during the training.
@@
    Players are bounded rational (Simon, 1955; Kahneman & Tversky, 1982; Selten, 1990). According to the concept of the bounded rationality, decision makers often struggle to find a global optimum due to their limited information, time, and computational resources. As a result, they remain at a nearby local optimum. Experimental studies (Nagel, 1995; Coricelli & Nagel, 2009) also support this claim by showing that players in games do not act rationally and their behavior is bounded. From a game theory perspective (Flåm, 1998; Chen et al., 2011; Ratliff et al., 2016), this implies that players tend to reach the local Nash equilibrium (Alos-Ferrer & Ania, 2001) and rarely shift to the different local optimal strategy.
Under these assumptions, the non-cheater uses the policy $\pi_{n}:\Omega\times\mathcal{A}\rightarrow[0,1]$, the cheater uses the policy $\pi_{c}:\mathcal{S}\times\Omega\times\mathcal{A}\rightarrow[0,1]$, and the cheat detector uses the classifier $D:\mathcal{T}\rightarrow[0,1]$, where $\mathcal{T}$ denotes the set of all possible trajectories. Since the non-cheater follows the local optimal policy under partial observability, $\pi_{n}$ can be obtained by solving a reward maximization problem $\max_{\pi_{n}}J(\pi_{n})$ with policy optimization algorithms such as A2C (Mnih et al., 2016) and PPO (Schulman et al., 2017), where $J(\pi_{n})$ denotes the expected return under policy $\pi_{n}$. Note that it is guaranteed to reach the local optimum when using temporal difference (Maei et al., 2009), deep Q-learning (Fan et al., 2020), and actor-critic methods (Holzleitner et al., 2021; Tian et al., 2023).
Next, we formulate the dynamic interaction between the ESP cheater and the detector as an adversarial game. In this setup, the detector aims to improve its classification performance by analyzing behavior patterns exhibited by the agents, while the cheater continually adapts its strategy to evade detection and maximize in-game rewards. Formally, the detector updates its classifier $D$ to distinguish between trajectories induced by the cheater policy $\pi_{c}$ and the non-cheater policy $\pi_{n}$, by minimizing the binary cross-entropy loss over samples $\tau_{c}\sim\pi_{c}$ and $\tau_{n}\sim\pi_{n}$. Simultaneously, the cheater optimizes its policy $\pi_{c}$ to maximize the expected return $J(\pi_{c})$ while minimizing $D(\tau_{c})$, thereby reducing the classifier’s confidence in identifying cheating behaviors. It is formalized as a minimax game:
@@
Note that Daskalakis & Panageas (2018); Adolphs et al. (2019); Mazumdar et al. (2020) theoretically proved that applying GDA to a minimax problem yields a local Nash equilibrium.
In practice, $\log D$ in equation 2 may not provide sufficiently strong gradients for effective optimization of $\pi_{c}$, as discussed in Goodfellow et al. (2020). When the classifier becomes confident in its predictions, $\log D(\tau)$ tends to saturate, thereby diminishing its influence on the objective. To alleviate this issue, we adopt $\log(1-D(\tau))$ in place of $\log D(\tau)$, as suggested in Goodfellow et al. (2020). It yields the following surrogate objective:
@@
To solve the equation 4, we apply reward shaping (Ng et al., 1999) with the shaped reward $r_{t}^{\prime}$:
@@
To reflect this behavior, we model the cheater policy $\pi_{c}$ as an interpolation of the non-cheater policy $\pi_{n}$ and the pure cheater policy $\pi_{c}^{(p)}:\mathcal{S}\times\Omega\times\mathcal{A}\rightarrow[0,1]$:
@@ 4.1 Experiment Setup
We employ three main metrics to evaluate the performances of the cheater and the detector: Average precision (AP), AUROC, and Average reward. AP and AUROC evaluate the detector’s ability to distinguish cheater trajectories from non-cheater ones, while the average reward measures the effectiveness of the policy in the game environment. Lower AP and AUROC indicate that the cheater successfully deceives the detector. In addition, we report Average trajectory length to provide further insights into the agent behavior. Under the non-adversarial setting, non-cheaters tend to produce longer trajectories than cheaters, since they must actively explore the environment to search for items or check the next card, leading to longer episode length. Lastly, we use Relative reward to visualize the reward changes over detectability. It is defined as (rewardnon-cheater’s reward)/(pure cheater’s rewardnon-cheater’s reward)($\text${reward}-$\text${non-cheater's reward})/($\text${pure cheater's reward}-$\text${non-cheater's reward}), to represent the normalized reward gain of the adversarial cheater compared to the pure cheater.
@@
We use a CNN-based actor-critic architecture for the policy network and a CNN-based classifier for the detector network. Both the cheater and non-cheater policies are trained using PPO algorithm (Schulman et al., 2017). Before conducting adversarial training, we construct the trajectory dataset for each policy and then pretrain the detector. The pretraining results are summarized in Tab. 1. For adversarial training, we finetune the cheater policy and the detector network based on Alg. 1. More details can be found in Appendix B.
@@ 5 Conclusions and Future Works
In this paper, we present the ESP simulation framework, which models ESP cheaters, non-cheaters, and cheat detectors with their adversarial relationships. Experimental results demonstrate that our framework can effectively simulate adaptive cheater behaviors that balance reward optimization and evade detection. We show that the fixed detector can be easily exploited by adaptive cheaters, emphasizing the need for continuously updated detection mechanisms. We further analyze that the cheater can outperform the non-cheater in terms of average reward, while remaining undetected by a trajectory-based detector. These results suggest that detecting adaptive cheaters may require more sophisticated approaches, such as utilizing behavior patterns observed across multiple episodes.

### v2

@@ 1 Introduction
Cheating is a persistent issue in digital games that significantly violates the fairness and degrades the user experience. If cheating is not properly addressed, it may drive players away from the game, ultimately leading to reduced revenue for game developers. It has been reported to cause an estimated loss of \$29 billion and to drive away 78% of gamers for a year (irdeto).
One of the most representative forms of cheating is Extra-Sensory Perception (ESP), which allows cheaters to access hidden game information. A common example is the use of wallhacks in first-person shooter (FPS) games: while normal players cannot see opponents behind walls, cheaters can, enabling them to avoid unexpected attacks and easily eliminate enemies. These unfair advantages severely disrupt the balance of the game, especially in settings where players must make decisions and formulate strategies based on incomplete information. ESP cheats are notoriously hard to prevent due to their passive nature: they do not modify game files or client-side data. In addition, their usage is often indistinguishable from normal gameplay when observed from a third-person perspective, making them especially challenging to detect through in-game monitoring or user reports (pubg).
@@
Next, cheaters continuously evolve their strategies to evade detection (tao2018nguard; jonnalagadda2021robust). When a cheat detector is developed and deployed, cheaters adapt by limiting their cheat usage just enough to avoid detection. For example, if a simple threshold-based detection algorithm is used, cheaters will intentionally operate just below the threshold to remain undetected. As a result, the statistical characteristics of past cheater data differ from those of recent data. Since only recent data is useful for developing effective detectors, it becomes difficult to collect a large amount of high-quality training and test data.
@@
To verify the effectiveness of the framework, we implement Gridworld and Blackjack environments that can simulate both fully observable and partially observable agents. These simplified single-agent environments provide interpretable behavioral indicators, such as exploration patterns and trajectory length, which allow us to quantitatively and qualitatively analyze the behaviors of cheaters under controlled conditions. Although these environments are abstracted, they capture fundamental aspects such as spatial exploration and reward-driven behavior that also appear in more complex game settings. In particular, even in FPS games, player movements and interactions can often be represented as grid-like trajectories on a minimap, making the Gridworld environment a meaningful abstraction for studying such dynamics. While our long-term goal is to extend the framework to complex multi-player environments such as FPS games, these experiments serve as a foundational baseline toward understanding the fundamental co-adaptation mechanism between cheaters and detectors, before extending the analysis to complex multi-agent environments. We also utilize various detector designs based on the trajectory, trajectory length, and reward. With these diverse configurations, we conduct extensive experiments that demonstrate the effectiveness of our framework in modeling adaptive cheaters. Our code is available at https://github.com/yunik1004/esp-simulator.
@@ 2 Related Works
Several studies (alayed2013behavioral; tao2018nguard; jonnalagadda2021robust; pinto2021deep; kanervisto2022gan; zhang2024identify) have explored cheat detection in video games, proposing a variety of approaches for identifying different types of malicious behaviors. alayed2013behavioral used a support vector machine (SVM) (suthaharan2016support) to detect aimbots based on server-side logs from online FPS games. tao2018nguard proposed NGUARD, a bot detection framework for multiplayer online role-playing games (MMORPGs), which combines supervised and unsupervised learning to detect known bots and discover previously unseen ones. jonnalagadda2021robust introduced a vision-based deep neural network (DNN) detector to identify illicit on-screen overlays in Counter-Strike: Global Offensive (CS:GO) (valve2012csgo). In addition, they applied the interval bound propagation method (gowal2018effectiveness) to defend against adversarial attacks targeting the detection system. pinto2021deep formulated sequences of keyboard and mouse events as a multivariate time series, and employed a convolutional neural network (CNN) (o2015introduction) to detect triggerbots and aimbots in CS:GO. kanervisto2022gan developed an aimbot that mimics human behavior using a generative adversarial network (GAN) (goodfellow2020generative), and evaluated the performance of multiple detection methods against it. zhang2024identify proposed HAWK, an anti-cheat framework for CS:GO, which leverages machine learning to mimic the decision process of human experts in detecting wallhacks and aimbots. Despite recent advances, most approaches assume that cheaters behave in fixed, non-adaptive ways. It overlooks the fact that real-world cheaters often adjust their strategies to avoid detection, highlighting the need for simulation frameworks that can model such adaptive behaviors.
@@
Adversarial learning traces its roots to the GAN (goodfellow2020generative), where a generator and a discriminator engage in a minimax game to match data distributions. ho2016_gail later carried this game-theoretic structure over to reinforcement learning with generative adversarial imitation learning (GAIL), which trains a policy to fool a discriminator that separates expert and learner trajectories, thereby reproducing the expert’s state–action occupancy without an explicit reward-recovery step. For robustness, pinto2017_rarl proposed robust adversarial reinforcement learning (RARL), where an adversary injects disturbance forces into the simulator so that the resulting policy remains stable under variations in friction, mass, and other model mismatches. zhang2020robust proposed state-adversarial Markov decision process (SA-MDP) framework to analyze the observation robustness of RL algorithms and introduced a policy regularization technique to enhance their robustness. sun2022who established a theoretical understanding of the optimality of adversarial attacks from the perspective of policy perturbations. zhang2021robust proposed alternating training with learned adversaries (ATLA), a framework that improves an agent’s robustness under perturbed observations by jointly training the adversary and the RL agent. franzmeyer2024illusory introduced the illusory attack, an information-theoretic adversarial framework that ensures statistically grounded detectability while maintaining attack effectiveness against both human and AI agents. dennis2020_paired presented protagonist–antagonist induced regret environment design (PAIRED), training an environment designer and an antagonist demonstrator to build a curriculum of tasks that are solvable yet currently unsolved by the protagonist, thereby boosting zero-shot transfer. The adversarial paradigm has even reached the language domain: the self-playing adversarial language game (ALG) (cheng2025_spag) places two large language models in a hide-and-seek game with taboo tokens, iteratively sharpening their reasoning strategies across benchmarks. Although these lines of work pursue diverse goals in imitation (ho2016_gail), robustness (pinto2017_rarl; zhang2020robust; zhang2021robust; sun2022who; franzmeyer2024illusory) to observation or policy perturbations, curriculum-driven generalization (dennis2020_paired), and domain expansion (cheng2025_spag), they all leverage adversarial interaction to enhance policy capability. In contrast, our framework jointly trains an ESP cheater and a trajectory-level detector in a partially observable setting where the detector provides the detection probability as a negative reward. This design captures the dynamic evolution of cheating behaviors that balance reward optimization and detection evasion.
Although the overall formulation of this work is similar to the minimax structure of GAN or GAIL, the modeling purpose is fundamentally different. GAN and GAIL aim to imitate the data distribution or policies, focusing on designing highly expressive generators and improving distribution-matching performance. In contrast, the cheater in our framework does not necessarily imitate the non-cheater’s policy. It instead learns new behaviors that balance the trade-off between detection avoidance and reward maximization. While imitation may help evade detection, it cannot achieve the cheater’s primary goal, which is obtaining unfair advantages through higher rewards. Another key difference of our study is that we treat the detector’s performance as an equally important subject of analysis. In conventional GAN or GAIL frameworks, the discriminator serves merely as an auxiliary tool for generator training, and its own performance is rarely analyzed. In contrast, our framework considers the detector as an active player. This approach reflects the unique characteristics of the cheat detection domain and provides a clear conceptual distinction from GAN and GAIL.
@@ 3 ESP Simulation Framework
We model the game environment as a partially observable Markov decision process (POMDP) (aastrom1965optimal; kaelbling1998planning), defined by the tuple $\langle\mathcal{S},\mathcal{A},T,r,\Omega,O,\gamma\rangle$, where $\mathcal{S}$ is the set of environment states, $\mathcal{A}$ is the set of actions, $T$ denotes the state transition probability distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\Omega$ is the set of observations available to the non-cheater player, $O$ is the observation function, defining the conditional observation probabilities, and $\gamma\in[0,1]$ is a discount factor. Consistent with previous works related to the trajectory feedback (liu2019sequence; efroni2021reinforcement), we assume $\gamma=1$ throughout this paper.
@@
    Given a trajectory, a sequence of states and actions $\tau=(s_{0},a_{0},s_{1},a_{1},\cdots,s_{t},a_{t})$, the detector estimates the probability that it was generated by the cheater. During the inference, the detector receives no information about the ground-truth label of the trajectory. On the other hand, it is allowed to access the ground-truth label during the training.
@@
    Players are bounded rational (simon1955behavioral; kahneman1982psychology; selten1990bounded). According to the concept of the bounded rationality, decision makers often struggle to find a global optimum due to their limited information, time, and computational resources. As a result, they remain at a nearby local optimum. Experimental studies (nagel1995unraveling; coricelli2009neural) also support this claim by showing that players in games do not act rationally and their behavior is bounded. From a game theory perspective (flaam1998restricted; chen2011community; ratliff2016characterization), this implies that players tend to reach the local Nash equilibrium (alos2001local) and rarely shift to the different local optimal strategy.
Under these assumptions, both the cheater and the non-cheater are using local optimal policies. It can produce a large performance gap between the two players, making it easier for the detector to distinguish them than in real-world scenarios. In other words, this setup represents an upper bound on the detector’s performance.
We now define the following components: the non-cheater’s policy $\pi_{n}:\Omega\times\mathcal{A}\rightarrow[0,1]$, the cheater’s policy $\pi_{c}:\mathcal{S}\times\Omega\times\mathcal{A}\rightarrow[0,1]$, and the cheat detector’s classifier $D:\mathcal{T}\rightarrow[0,1]$, where $\mathcal{T}$ denotes the set of all possible trajectories. Since the non-cheater follows the local optimal policy under partial observability, $\pi_{n}$ can be obtained by solving a reward maximization problem $\max_{\pi_{n}}J(\pi_{n})$ with policy optimization algorithms such as A2C (mnih2016asynchronous) and PPO (schulman2017proximal), where $J(\pi_{n})$ denotes the expected return under policy $\pi_{n}$. Note that it is guaranteed to reach the local optimum when using temporal difference (maei2009convergent), deep Q-learning (fan2020theoretical), and actor-critic methods (holzleitner2021convergence; tian2023convergence).
We then formulate the dynamic interaction between the ESP cheater and the detector as an adversarial game. In this setup, the detector aims to improve its classification performance by analyzing behavior patterns exhibited by the agents, while the cheater continually adapts its strategy to evade detection and maximize in-game rewards. Formally, the detector updates its classifier $D$ to distinguish between trajectories induced by the cheater policy $\pi_{c}$ and the non-cheater policy $\pi_{n}$, by minimizing the binary cross-entropy loss over samples $\tau_{c}\sim\pi_{c}$ and $\tau_{n}\sim\pi_{n}$. Simultaneously, the cheater optimizes its policy $\pi_{c}$ to maximize the expected return $J(\pi_{c})$ while minimizing $D(\tau_{c})$, thereby reducing the classifier’s confidence in identifying cheating behaviors. It is formalized as a minimax game:
@@
Note that daskalakis2018limit; adolphs2019local; mazumdar2020gradient theoretically proved that applying GDA to a minimax problem yields a local Nash equilibrium.
In practice, $\log D$ in equation 2 may not provide sufficiently strong gradients for effective optimization of $\pi_{c}$, as discussed in goodfellow2020generative. When the classifier becomes confident in its predictions, $\log D(\tau)$ tends to saturate, thereby diminishing its influence on the objective. To alleviate this issue, we adopt $\log(1-D(\tau))$ in place of $\log D(\tau)$, as suggested in goodfellow2020generative. It yields the following surrogate objective:
@@
To solve the equation 4, we apply reward shaping (ng1999policy) with the shaped reward $r_{t}^{\prime}$:
@@
To reflect this behavior, we model the cheater policy $\pi_{c}$ as an interpolation of the non-cheater policy $\pi_{n}$ and the pure cheater policy $\pi_{c}^{(p)}:\mathcal{S}\times\Omega\times\mathcal{A}\rightarrow[0,1]$, inspired by the mixture-of-experts architecture (jacobs1991adaptive; shazeer2017outrageously):
@@ 4.1 Experiment Setup
We employ three main metrics to evaluate the performances of the cheater and the detector: Average Precision (AP), Area Under the Receiver Operating Characteristic curve (AUROC), and Average reward. AP measures the area under the precision–recall curve, and AUROC measures the area under the ROC curve representing the trade-off between true and false positive rates. A higher AUROC reflects a stronger ability to identify cheaters and avoid false alarms. AP and AUROC evaluate the detector’s ability to distinguish cheater trajectories from non-cheater ones, while the average reward measures the effectiveness of the policy in the game environment. Lower AP and AUROC indicate that the cheater successfully deceives the detector. In addition, we report Average trajectory length to provide further insights into the agent behavior. Under the non-adversarial setting, non-cheaters tend to produce longer trajectories than cheaters, since they must actively explore the environment to search for items or check the next card, leading to longer episode length. Lastly, we use Relative reward to visualize the reward changes over detectability. It is defined as (rewardnon-cheater’s reward)/(pure cheater’s rewardnon-cheater’s reward)($\text${reward}-$\text${non-cheater's reward})/($\text${pure cheater's reward}-$\text${non-cheater's reward}), to represent the normalized reward gain of the adversarial cheater compared to the pure cheater.
@@
We use a CNN-based actor-critic architecture for the policy network and a CNN-based classifier for the detector network. Both the cheater and non-cheater policies are trained using PPO algorithm (schulman2017proximal). Before conducting adversarial training, we construct the trajectory dataset for each policy and then pretrain the detector. The pretraining results are summarized in Tab. 1. For adversarial training, we finetune the cheater policy and the detector network based on Alg. 1. More details can be found in Appendix B.
@@ 5 Conclusions and Future Works
In this paper, we present the ESP simulation framework, which models ESP cheaters, non-cheaters, and cheat detectors with their adversarial relationships. Experimental results demonstrate that our framework can effectively simulate adaptive cheater behaviors that balance reward optimization and evade detection. We show that the fixed detector can be easily exploited by adaptive cheaters, emphasizing the need for continuously updated detection mechanisms. Although the setup is disadvantageous to the cheater, we further analyze that the cheater can outperform the non-cheater in terms of average reward, while remaining undetected by a trajectory-based detector. These results suggest that detecting adaptive cheaters may require more sophisticated approaches, such as utilizing behavior patterns observed across multiple episodes.